When people visit a certain large shop, on average 34% of them do not buy anything, 53% spend less 

than $50 and 13% spend at least $50.

(i) 15 people visiting the shop are chosen at random. Calculate the probability that at least 14 of 

them buy something.

(ii) n people visiting the shop are chosen at random. The probability that none of them spends at 

least $50 is less than 0.04. Find the smallest possible value of n.



1
Expert's answer
2021-10-04T16:29:26-0400

Solution:

"n=15,q=0.34,p=0.66"

(i) "P(X\ge 14)=P(X=14)+P(X=15)"

"=^{15}C_{14}(0.66)^{14}(0.34)^1+^{15}C_{15}(0.66)^{15}(0.34)^0
\\=0.0171"

(ii) "p=0.13,q=0.87"

"P(X=0)<0.04"

"\Rightarrow ^nC_0(0.87)^n(0.13)^0<0.04"

"\Rightarrow (0.87)^n<0.04"

Using hit and trial, put n=23

"LHS=(0.87)^{23}=0.041>0.04"

Put n=24.

"LHS=(0.87)^{24}=0.035<0.04"

Thus, required value of n=24.